5) ʃ (1-sinx /1-cosx)ex dx

6) ʃ0π/4 cosθ +isinθ / 2cos2θ dθ


1
Expert's answer
2021-09-14T06:09:38-0400

5). "\int ({1-sin(x)\over1-cos(x)})e^xdx"

Rewrite to get:

"\int ({sin(x)-1\over cos(x)-1})e^xdx"

Rewriting using trigonometric identities:

"\int ({csc^2({x\over 2})e^x\over 2}-cot({x\over 2})e^x)dx"

Applying linearity we get:

"{1\over 2}\int csc^2({x\over 2})e^xdx-\int cot({x\over 2})e^xdx"

Applying integration by parts for "\int cot({x\over 2})e^xdx"

Let "u=cot({x\over 2}) \implies du=-{csc^2({x\over 2})e^x\over 2}dx"

Let "dv=e^xdx \implies v=e^x"

"\int udv=uv-\int vdu"

"=cot({x\over 2})e^x-\int -e^x{csc^2({x\over 2})e^x\over 2}dx"

"=cot({x\over 2})e^x+\int e^x{csc^2({x\over 2})e^x\over 2}dx"

"=cot({x\over 2})e^x+{1\over 2}\int csc^2({x\over 2})e^xdx"

"\implies \int cot({x\over 2})e^xdx=cot({x\over 2})e^x+{1\over 2}\int csc^2({x\over 2})e^xdx"

"\implies {1\over 2}\int csc^2({x\over 2})e^xdx-\int cot({x\over 2})e^xdx={1\over 2}\int csc^2({x\over 2})e^xdx-[cot({x\over 2})e^x+{1\over 2}\int csc^2({x\over 2})e^xdx]"

"={1\over 2}\int csc^2({x\over 2})e^xdx-cot({x\over 2})e^x-{1\over 2}\int csc^2({x\over 2})e^xdx"

The integral "{1\over 2}\int csc^2({x\over 2})e^xdx" cancels,hence we get

"=-cot({x\over 2})e^x+C"

"\therefore \int ({1-sin(x)\over1-cos(x)})e^xdx=-cot({x\over 2})e^x+C"


6). "\int_0^{\pi\over 4} {cos(\theta)+sin(\theta)\over 2cos^2(\theta)}d\theta"

Rewrite to get:

"=\int_0^{\pi\over 4} ({cos(\theta)\over 2cos^2(\theta)}+{sin(\theta)\over 2cos^2(\theta)})d\theta"

Applying linearity we get:

"={1\over 2}\int_0^{\pi\over 4}{cos(\theta)\over cos^2(\theta)}d\theta+{1\over 2}\int_0^{\pi\over 4}{sin(\theta)\over cos^2(\theta)})d\theta"

"={1\over 2}\int_0^{\pi\over 4}{1\over cos(\theta)}d\theta+{1\over 2}\int_0^{\pi\over 4}{sin(\theta)\over cos^2(\theta)})d\theta"

"={1\over 2}[\int_0^{\pi\over 4}{1\over cos(\theta)}d\theta+\int_0^{\pi\over 4}{sin(\theta)\over cos^2(\theta)})d\theta]"

Applying integration by substitution for "\int_0^{\pi\over 4}{sin(\theta)\over cos^2(\theta)})d\theta" we have

Let "u=cos(\theta) \implies du=-sin(\theta)d\theta"

"\implies \int_0^{\pi\over 4}{sin(\theta)\over cos^2(\theta)})d\theta=-\int_0^{\pi\over 4}{du\over u^2}=-\int_0^{\pi\over 4}u^{-2}du=[-u^{-1}]_0^{\pi\over 4}=[-{1\over cos(\theta)}]_0^{\pi\over 4}=-[{1\over cos({\pi\over 4})}-{1\over cos(0)}]={1\over cos(0)}-{1\over cos({\pi\over 4})}"

"\therefore \int_0^{\pi\over 4}{sin(\theta)\over cos^2(\theta)})d\theta={1\over cos(0)}-{1\over cos({\pi\over 4})}=1-1.41421356237309=-0.41421356237309"

Now, let's find "\int_0^{\pi\over 4}{1\over cos(\theta)}d\theta"

"{1\over cos(\theta)}=sec(\theta)"

"\implies \int_0^{\pi\over 4}{1\over cos(\theta)}d\theta= \int_0^{\pi\over 4}sec(\theta)d\theta"

Expand fraction by "(tan(\theta)+sec(\theta))"

"\int_0^{\pi\over 4}sec(\theta)d\theta= \int_0^{\pi\over 4}sec(\theta){tan(\theta)+sec(\theta)\over tan(\theta)+sec(\theta)}d\theta"

"= \int_0^{\pi\over 4}{sec(\theta)tan(\theta)+sec^2(\theta)\over tan(\theta)+sec(\theta)}d\theta"

Applying integration by substitution for "\int_0^{\pi\over 4}{sec(\theta)tan(\theta)+sec^2(\theta)\over tan(\theta)+sec(\theta)}d\theta"

Let "u=tan(\theta)+sec(\theta) \implies du=[sec(\theta)tan(\theta)+sec^2(\theta)]d\theta"

"\implies d\theta ={du\over sec(\theta)tan(\theta)+sec^2(\theta)}"

"\implies\int_0^{\pi\over 4}{sec(\theta)tan(\theta)+sec^2(\theta)\over tan(\theta)+sec(\theta)}d\theta=\int_0^{\pi\over 4}{sec(\theta)tan(\theta)+sec^2(\theta)\over u}{du\over sec(\theta)tan(\theta)+sec^2(\theta)}"

The terms "sec(\theta)tan(\theta)+sec^2(\theta)" cancels hence we get

"=\int_0^{\pi\over 4}{du\over u}=[ln(u)]_0^{\pi\over 4}=[ln(tan(\theta)+sec(\theta))]_0^{\pi\over 4}=[ln(tan({\pi\over 4})+sec({\pi\over 4}))]-[ln(tan(0)+sec(0))]"

"tan({\pi\over 4})=1"

"sec({\pi\over 4})={1\over cos({\pi\over 4})}=1.41421356237309"

"tan(0)=0"

"sec(0)={1\over cos(0)}=1"

"\int_0^{\pi\over 4}{du\over u}=[ln(1+1.41421356237309)]-[ln(0+1)]"

"=[ln(2.41421356237309)]-[ln(1)]=0.88137358701954"

"\implies \int_0^{\pi\over 4} {cos(\theta)+sin(\theta)\over 2cos^2(\theta)}d\theta={1\over 2}[0.88137358701954-0.41421356237309]=0.23358001232323"